Blast it! said David Wilson, president of Teledex Company. Weve just lost the bid on the Koopers job by $3,000. It seemsRequired: 1. Using the companys plantwide approach: a. Compute the plantwide predetermined rate for the current year. b. Det
Answer #1

1a) Plantwide overhead rate = 903000/645000 = 140% of labor cost

1b) Amount of overhead applied

Fabricating Machining Assembly Total
Manufacturing overhead applied 5800*1.4 = 8120 800*1.4 = 1120 7700*1.4 = 10780 20020

2a) Departmental overhead rate

Fabricating = 376250/215000 = 175% of labor cost

Machining = 430000/107500 = 400% of labor cost

Assembly = 96750/322500 = 30% of labor cost

2b) Overhead applied

Fabricating Machining Assembly Total
Manufacturing overhead applied 5800*1.75 = 10150 800*4 = 3200 7700*30% = 2310 15660

3a) Bid price = 7900+14300+20020*1.5 = 63330

3b) Bid price = 7900+14300+15660*1.5 = 56790